# Non-unital noise in a superconducting quantum computer as a computational resource for reservoir computing

### v2

@@ Abstract
Despite theoretical promises, most machine learning algorithms fail on current quantum computers due to overwhelming noise. Nevertheless, many recurrent architectures are most efficient if they can process information by progressively reducing the correlation on inputs from earlier time steps. Among them, reservoir computing represents a major paradigm of artificial intelligence for processing time-dependent tasks thanks to the involvement of fading memory. Despite attempts to demonstrate quantum reservoir computing based on intentional perturbations of the network, whether the intrinsic noise of the quantum circuit can be straightforwardly exploited remained unknown. We show the feasibility of reservoir computing on a circuit of superconducting qubits based on the unavoidable dissipation typical of NISQ devices. We prove that noise modeled by a non-unital quantum channel ensures the functioning of a quantum echo state network. In particular, amplitude damping is responsible for drastically improving the short-term memory capacity and expressivity of the network, by simultaneously providing fading memory and richer dynamics. Our experimental results pave the way for the application of reservoir computing methods in non-fault-tolerant quantum computers.
@@ Introduction
The intrinsic dissipation and decoherence in noisy intermediate-scale quantum (NISQ) computers pose a major limitation in many machine-learning protocols [1], restricting the natural application of quantum computing in artificial intelligence [2, 3]. Nevertheless, dissipation may serve as a fully-fledged resource for quantum computation [4, 5, 6]. Our results show that the action of specific quantum noise in the hardware, namely noise modeled by non-unital quantum channels, significantly improves the short-term memory capacity and expressivity of a quantum network, proving the reliability of quantum reservoir computing as a valid computational architecture for memory-dependent tasks employing a gate-based quantum computer. We emphasize that an in-depth study of the beneficial computational effects of non-unital noise in quantum machine learning is only in its early stages [7], unveiling, for example, a promising application in avoiding barren plateaus in variational problems [8]. ng Reservoir computing is a well-established supervised machine learning algorithm that employs a fixed neural network – the reservoir, to process time-dependent information. In addition to initial digital and neuromorphic proposals, respectively, the echo state network [9] and the liquid state machine [10], reservoir computing has proven effective in a wide range of unconventional deployments [11]. Among several physical implementations [12], the exponentially large computational capacity of quantum systems has been harnessed for reservoir computing [13, 14], employing photonic circuits [15, 16, 17], bosonic oscillators [18, 19], fermionic systems [20, 21], neutral atoms [22] and spin networks [23, 24, 25, 26, 27, 28]. Focusing on superconducting quantum computers, we refer to Ref. [29, 30] for early implementations, more recently empowered by mid-circuit measurements [31]. Reservoir computing appears particularly suitable for leveraging quantum computing, as it does not require any training in the parameters of the quantum evolution. An essential feature essential for reservoir computing is the so-called fading memory [32, 33]. Namely, a reservoir network is most efficient if the information is processed by progressively reducing the emphasis on inputs from earlier time steps. Dissipation has recently been recognized as a powerful computational resource in reservoir architectures that employ spin dynamics [27]. In contrast, phase errors seem to have no utility for quantum computing, since they are theoretically associated with loss of quantum information. Concerning reservoir computing employing superconducting quantum computers, artificial manual reset of the information flow has proven to be a reliable resource for fading memory in memory-dependent tasks [30, 34]. Here, we step further by demonstrating how the intrinsic noise in superconducting quantum computers can be harnessed to spontaneously ensure the functioning of a quantum echo state network. We implement reservoir computing on a circuit of 7 superconducting qubits and we perform its emulation by including a realistic noise model. By testing our quantum echo state network on standard memory-dependent benchmarks, we recognize which type of noise benefits the performances. Thus, we show that noise modeled by a non-unital channel is needed to simultaneously guarantee memory capacity and accuracy in nonlinear time-dependent tasks. This confirms early results for memory-independent classification tasks [35]. We recall that a quantum channel is unital if it preserves the identity in the operator space. Remarkably, the quantum channel modeling the noise of a superconducting quantum computer falls within such a class of non-unital noise. As typically observed in recurrent networks [36, 37], we identify a critical regime tuned by noise intensity, in which several network capabilities, such as short-term memory capacity and expressivity, are maximized. In this respect, we suitably slowed the execution of the circuit accordingly to maximize the learning accuracy. Finally, the universality of a neural network architecture is responsible for its computational effectiveness, since it ensures that any input-output mapping can be approximated with arbitrary precision [38, 10, 39, 29, 18, 27, 40]. Here, we provide strong theoretical support to our empirical findings, by proving the universality of our gate-based echo state network under the effect of non-unital quantum channels. Our experiment paves the way for real-world application of quantum reservoir computing on noisy intermediate-scale and early fault-tolerant quantum computers.
@@ Results
In order to demonstrate the feasibility of quantum reservoir computing on a noisy gate-based quantum computer, we proceed by first defining the embodiment of reservoir computing over the qubits. Next, we emulate the response of the reservoir by involving different kinds of noise to confirm that – as predicted, only the non-unital noise guarantees non-trivial dynamics required for the learning. According to realistic noise models, being non-unital noise a relevant component in that of superconducting quantum computers, we, therefore, move to such hardware implementation to demonstrate experimentally the learning. As the amount of noise determines the accuracy of the learning, the experiment takes into account a tuning of the execution time in order to maximize the accuracy.
Figure 1: The architecture of quantum reservoir computing. a) The time-dependent information in the input is sampled to produce the input time series. The information is stored in the logical nodes described by the Pauli basis associated with the qubits. During the evolution of the reservoir, the information is extracted by mid-circuit measurements of the true nodes (in green) of the network. The reservoir signal is the outcome of the measurements. The only trained part of the system is the readout weights $\omega$ (red connections). b) The circuit is employed as the reservoir of the quantum echo state network. The input values are encoded in parametric unitary gates. Mid-circuit measurements are performed, to extract the reservoir signal online. After the execution, the optimal weights of the linear readout are computed through the pseudoinverse of the overall reservoir signal along the training interval, in order to minimize the distance from the target time series. c) The precise architecture of the gates in each unitary layer. The inputs are encoded in the composition of rotations along the $X$ (in blue) and $Z$ (in pink) axis of the Bloch sphere. Z Pauli operators are measured as true nodes. CNOT operators allow for the entanglement of the qubits. The qubits are left in the collapsed state after each measurement. In the work, we fix $N=7$ as the number of qubits. Finally, the reservoir signal, namely the expected value of each Z operator, is collected in the matrix $H$.
Problem setting
The task consists of defining a mapping – usually called filter, in the framework of reservoir computing, sending an input time series $u\in\mathcal{I}$ into another target time series $y\in\mathcal{O}$, through read-out operations of the information encoded in the reservoir. The schematic structure is represented in Fig. 1. Denoting $\mathcal{S}(\mathcal{H})$ the set of density operators $\rho(t)$ on the Hilbert space $\mathcal{H}$ that describes some quantum system, a quantum reservoir computer is described by the equations
@@
The tasks are selected among well-established benchmarks in the literature of reservoir computing, more specifically we consider NARMAn tasks, up to $n=8$, as described in the Methods. Both the emulator of the IBM-Osaka and the real quantum computer IBM-Brisbane have been exploited to perform the learning tasks. For what concerns the simulation of diverse models of quantum noise, their effect on the learning capabilities of the network is investigated by systematically tuning its intensity in numerical experiments. Throughout the whole Results section, the parameter that tunes noise intensity is referred to with $\gamma$.
@@
where the vector $\overline{z}^{i}_{t}=\left\{(z_{t})_{s}\right\}_{s=1,\dots,S}$ contains the outcomes of measurements for each repetition of the experiment. Then, we compute the mean correlation over the whole execution by averaging over $t$. Fig. 2 shows the mean correlation over time for increasing time windows $k$. As expected, higher noise intensity enhances the progressive loss of correlation. The loss of correlation appears comparable for different noise models. Thus, fading memory alone does not a priori justify the beneficial effect of amplitude damping on the short-term memory capacity of the network and its accuracy in learning tasks. From a theoretical point of view, fading memory is related to the contractivity of the evolution of the echo state network [39]. In our model, the contractivity in trace norm is ensured by the action of noise, as discussed in the Supplementary Material.
@@
Figure 2: Effect on different noises on learning. a) Decay over time in the mean correlation in subsequent measurements is enhanced by increasing amplitude damping intensity. Yellow lines indicate a higher intensity of the noise. The thickness of lines is proportional to the accuracy in NARMA5 prediction. b) From left to right, the blue area is the image of the quantum channel on the Bloch sphere. Black and blue dots are the projection of the sphere’s center, respectively, before and after the application of the channel. For clarity, simulations are for $\gamma=0.6$ and tenfold intensity of noise in the IBM_OSAKA backend. c) Accuracy in NARMA task and short-term memory capacity for a different number of measured qubits in the 7 qubits register. Yellow lines correspond to complete measurements. Running average on a window of ten values of $\gamma$ is performed for smoother visualization. d) Accuracy in NARMAn tasks with complete measurement of the register under the effect of increasing amplitude damping. Results for a swarm of 20 executions are plotted in transparency, mean values are in bold in the picture. Running average on a window of ten values of $\gamma$ is performed for smoother visualization. The optimal noise regime corresponds to $\gamma\sim 0.03$. e) Accuracy in reproducing NARMA tasks and memory capacity for increasing intensity of amplitude damping (in blue), phase damping (in grey), and depolarizing channel(in green). Only amplitude damping ensures the learning of the network.
@@
Figure 3: Recostruction of NARMA tasks on simulators and quantum hardware. a) The reconstruction of NARMAn tasks for, respectively, the noiseless execution of the circuit, the execution affected by optimal amplitude damping, and the realistic noise emulating the one in IBM_OSAKA quantum backend. Optimal values of noise are $\gamma=0.024,0.03,0.014$ for NARMA2, NARMA5 and NARMA8, respectively. The grey vertical line separates the training and test interval in the output sequence. The amplitude damping ensures the best learning of the system, which is slightly deteriorated under the composition of several noisy quantum channels representing the realistic noise. b) The experimental results of reconstruction of NARMA2, NARMA5 and NARMA8 (from left to right) tasks employing IBM_BRISBANE quantum backend. The experiment is repeated 5 times. The mean results are reported in the green line, with error bars indicating a $1\,\mathrm{s.d.}$ interval. The input length is reduced, in order to reduce the depth of the circuit to fit the technical limitations of the quantum hardware. Precisely, we sample 80 values for $u_{t}$. Reconstruction deteriorates for increasing the depth in memory required, indicating the necessity of the implementation of error mitigation techniques that allow for reducing the unital noise in the evolution.
Effect on learning of different simulated quantum noise models
@@
After confirming that any contractive noisy channel ensures the fading memory of the network, we investigate their beneficial effect on the predictive capabilities of the architecture using two standard benchmarks for recurrent networks, namely the measure of short-term memory capacity (MC) and the non-linear autoregressive moving average (NARMA) task, presented in detail in Subsection ”Benchmark for performance analysis”. The short-term memory capacity measures the ability of the network to reproduce online a delayed window of the time-dependent input, while the NARMA task requires approximating a non-linear functional that depends on a fixed amount of the recent information in the input time series. Three noise models are applied separately, namely amplitude damping, phase damping, and depolarizing. Among them, only the amplitude damping is non-unital Remarkably, we observe that the action of the only amplitude damping drastically enhances the memory capacity of the network. Consistently, the higher memory capacity is reflected in better accuracy in reproducing the time-dependent task. On the other hand, phase damping and depolarizing noise have detrimental effects on both memory capacity and accuracy. These results are summarized in Fig. 2, showing the normalized mean square error (NMSE) and MC for increasing the intensity of the three noises considered. The beneficial effect of amplitude damping is justified by its non-unital nature as a quantum channel. Indeed, the action of a non-unital channel enriches the dynamics of the reservoir allowing for a wider exploration of the phase space and, consequently, greater storage of information. We discuss more details about non-unital channels and theoretical insight into the role of unitality in quantum reservoir computing in the Supplementary Material. To anticipate the behavior on NISQ hardware, we repeated the experiment by applying a channel that replicates the noise model of the backends provided by IBM. A realistic noise model for current superconducting hardware consists of the application of the composition of quantum channels, including, but not restricted to, the three noise models considered in this work. Thus, recalling that quantum channels are non-expanding by definition [42], we may assume that the dynamics is strictly contractive since the composition of non-expanding channels and strictly contractive channels is in turn a contractive channel. Moreover, we confirmed that the realistic noise is non-unital by analyzing its effect on the Bloch sphere, as shown in Fig. 2. As expected, the network still shows a remarkable improvement in learning capabilities when compared with noiseless executions, even if the accuracy in reproducing tasks is slightly deteriorated compared to optimal pure amplitude damping. The results in the reconstruction of the NARMA task are reported in Fig. 3.

### v3

@@ Abstract
Despite theoretical promises, most machine learning algorithms fail on current quantum computers due to overwhelming noise. Indeed, noise may limit their ability to retain and manipulate information over time, which is especially problematic when dealing with sequential data. Nevertheless, many recurrent architectures are most efficient if they can process information by progressively reducing the correlation on inputs from earlier time steps. Among them, reservoir computing represents a major paradigm of artificial intelligence for processing time-dependent tasks thanks to the involvement of fading memory. Despite attempts to demonstrate quantum reservoir computing based on intentional perturbations of the network, why the intrinsic noise of the quantum circuit can be straightforwardly exploited remained unknown. We show the feasibility of reservoir computing on a circuit of superconducting qubits based on the dissipation typical of NISQ devices. We prove that noise modeled by a non-unital quantum channel ensures the functioning of a quantum echo state network. In particular, amplitude damping is responsible for drastically improving the short-term memory capacity and reconstruction capability of the network by simultaneously providing fading memory and richer dynamics. Indeed, the damping is capable of preventing the loss of information due to repeated measurements of the circuit. Our experimental results pave the way for the application of reservoir computing methods in non-fault-tolerant quantum computers.
@@ Introduction
The intrinsic dissipation and decoherence in noisy intermediate-scale quantum (NISQ) computers pose a major limitation in many machine-learning protocols [1], restricting natural applications of quantum computing in artificial intelligence [2, 3]. Nevertheless, dissipation may serve as a resource for quantum computation [4, 5, 6]. Our results show that specific quantum noise of the hardware, namely noise modeled by non-unital quantum channels, significantly improves the short-term memory capacity and expressivity of a quantum network, proving the reliability of quantum reservoir computing as a valid computational architecture for memory-dependent tasks employing a gate-based quantum computer. We emphasize that an in-depth study of the beneficial effects of non-unital noise in quantum machine learning is only in its early stages [7], unveiling, for example, a promising application in avoiding barren plateaus in variational problems [8]. Reservoir computing is a well-established supervised machine learning algorithm that employs a fixed neural network – the reservoir, to process time-dependent information. In addition to initial digital and neuromorphic proposals, respectively, the echo state network [9] and the liquid state machine [10], reservoir computing has proven effective in a wide range of unconventional deployments [11]. Among several physical implementations [12], the exponentially large computational capacity of quantum systems has been harnessed for reservoir computing [13, 14], employing photonic circuits [15, 16, 17], bosonic oscillators [18, 19], fermionic systems [20, 21], neutral atoms [22] and spin networks [23, 24, 25, 26, 27, 28]. Focusing on superconducting quantum computers, we refer to Ref. [29, 30] for early implementations, more recently empowered by mid-circuit measurements [31]. Reservoir computing appears particularly suitable for leveraging the high-dimensional computational space associated with a quantum system [32], as it does not require any training in the parameters of the quantum evolution, thus avoiding typical issues associated with the training of a variational quantum circuit [33, 34]. An essential feature for reservoir computing is the so-called fading memory [35, 36]. Namely, a reservoir network is most efficient when dealing with the approximation of functional with past dependence if the information is processed by progressively reducing the emphasis on inputs from earlier time steps. Dissipation has recently been recognized as a powerful computational resource in reservoir architectures that employ spin dynamics [27]. In contrast, phase errors seem to have no utility for quantum computing, since they are theoretically associated with the loss of quantum information. Concerning reservoir computing employing superconducting quantum computers, artificial manual reset of the information flow has proven to be a reliable resource for fading memory in memory-dependent tasks [30, 37, 38]. In this work, we advance previous approaches by demonstrating why the inherent noise in superconducting quantum processors can be exploited to drive the dynamics required for an effective quantum echo state network. In particular, we show that the damping is essential to guarantee the effective functioning of a gate-based echo state network that operates relying on mid-circuit measurements. Indeed, dissipation can prevent the loss of information due to repeated measurements of the quantum circuit, which steers the internal state of the reservoir toward the maximally mixed state, making the encoding of new input values impossible. Thus, we provide strong theoretical support to our results and to previous empirical findings, which already highlighted the benefit of using amplitude damping in quantum reservoir computing [31, 39]. We implement reservoir computing on a circuit of 7 superconducting qubits, and we perform its emulation by including a realistic noise model. By testing our quantum echo state network on standard memory-dependent benchmarks, we recognize which type of noise benefits the performance. Thus, we show that noise modeled by a non-unital channel is needed to simultaneously guarantee memory capacity and accuracy in nonlinear time-dependent tasks. This confirms early results for memory-independent classification tasks [39]. We recall that a quantum channel is unital if it preserves the identity in the operator space. Remarkably, the quantum channel modeling the noise of a superconducting quantum computer falls within such a class of non-unital noise. As typically observed in recurrent networks [40, 41], we identify a critical regime tuned by noise intensity, in which several network capabilities, such as short-term memory capacity and expressivity, are maximized. In this respect, we suitably slowed the execution of the circuit accordingly to maximize the learning accuracy. Finally, as a further theoretical result, we prove the universality of our gate-based echo state network under the effect of non-unital quantum channels [42], enlarging the class of known universal quantum reservoir computers [18, 27, 29]. We recall that the universality of a neural network architecture is responsible for its computational effectiveness, since it ensures that any input-output mapping can be approximated with arbitrary precision [43, 10, 44]. Our experiment paves the way for real-world application of quantum reservoir computing on noisy intermediate-scale and early fault-tolerant quantum computers.
@@ Results
Problem settings
The task of a reservoir computer for the processing of temporal data consists of defining a mapping – usually called filter, in the framework of reservoir computing, sending an input time series $u\in\mathcal{I}$ into another target time series $y\in\mathcal{O}$, through read-out operations of the information encoded in the reservoir. The schematic structure is represented in Fig. 1. Denoting $\mathcal{S}(\mathcal{H})$ the set of density operators $\rho_{t}$ on the Hilbert space $\mathcal{H}$ that describes some quantum system, a quantum reservoir computer is described by the equations
@@
To demonstrate the feasibility of quantum reservoir computing on a noisy gate-based quantum computer, we proceed by first proving that some properties of the quantum channel $\mathcal{T}$, specifically non-unitality and contracivtit, satisfied by amplitude damping, ensure the effective functioning of a gate-based echo state network relying on mid-circuit measurements, thus furnishing strong theoretical support to the known results [31, 39]. Then, we show the positive impact of amplitude damping with respect to other noisy channels by employing well-established benchmark tasks in the literature of reservoir computing. More specifically, we consider NARMA-n tasks, up to $n=8$, as described in the Methods. Given that realistic noise models indicate non-unital noise as a significant feature of superconducting quantum computers, we proceed with a hardware implementation on an IBM quantum computer to experimentally validate the learning capabilities.
Figure 1: The architecture of quantum reservoir computing. a) The time-dependent information in the input is sampled to produce the input time series. The information is stored in the logical nodes described by the Pauli basis associated with the qubits. During the evolution of the reservoir, the information is extracted by mid-circuit measurements of the true nodes (in green) of the network. The reservoir signal is the outcome of the measurements. The only trained part of the system is the readout weights $\omega$ (red connections). b) The circuit is employed as the reservoir of the quantum echo state network. The input values are encoded in parametric unitary gates. Mid-circuit measurements are performed to extract the reservoir signal online. After the execution, the optimal weights of the linear readout are computed through the pseudoinverse of the overall reservoir signal along the training interval, in order to minimize the distance from the target time series. c) The precise architecture of the gates in each unitary layer. The inputs are encoded in the composition of rotations along the $X$ (in blue) and $Z$ (in pink) axis of the Bloch sphere. Z Pauli operators are measured as true nodes. CNOT operators allow for the entanglement of the qubits. The qubits are left in the collapsed state after each measurement. In the work, we fix $N=7$ as the number of qubits. Finally, the reservoir signal, namely the expected value of each Z operator, is collected in the matrix $H$.
Figure 2: Information retention induced by damping. a-b) Mean trace distance of the internal state of the reservoir after input injection for, respectively, (a.) increasing distance between inputs and (b.) increasing trace distance from $\hat{\mathcal{I}}$, recalling that this is bounded by $1-\frac{1}{2^{N_{a}}}$. In the first case, the average is computed over 100 initial states $\rho_{0}$, with distance from $\hat{\mathcal{I}}$ sampled from a uniform distribution $U(0,\frac{1}{2^{N_{a}}})$. In the second case, the average is computed over 100 uniformly sampled input values. c) Trace distance of the circuit from the $\hat{\mathcal{I}}$ along 40 time steps of input encoding and mid-circuit measurements, averaged over 10 repetitions with random inputs. If the dynamics is affected by amplitude damping (full line), the state remains deviated from $\hat{\mathcal{I}}$ during the evolution d) A sketch of the evolution of the internal state of a 1-qubit circuit reservoir during three steps of input encoding and midcircuit measurements. The unitary evolution does not modify the distance from the origin of the Bloch sphere, while the measurement collapses the state, reducing the trace distance from the maximally mixed state. The damping prevents the information loss by steering the state towards |$\ket{0}$.
Quantum reservoir computing with mid-circuit measurements
An approach to gate-based quantum reservoir computing, also pursued in this work, involves dealing with a fully accessible reservoir of $N_{a}$ qubits [46, 31], without any reset of the qubits, and thus exploiting quantum trajectories as a memory retention mechanism. Specifically, the input encoding is performed through parametric unitary matrices, and the information is decoded by measuring the qubits online along the execution of the circuit by exploiting mid-circuit measurements. After each measurement, the qubits are let evolve starting from the collapsed state. Thus, information retention relies on the statistics of the observables of the accessible qubits only, without the need for a hidden reservoir, as described in the Methods. Nonetheless, it is well known that repeated measurements on a quantum system progressively induce information loss. This hinders the computational capabilities of the reservoir computer, as we discuss later in this Section. Our work shows how the controlled damping of the qubits can overcome this issue, making the algorithm stable over time intervals significantly longer than the typical information survival time in the measured system. In particular, this work proves that the effect of damping compensates the quantum information loss due to mid-circuit measurements, thereby ensuring the persistence of the algorithm. We therefore explain also previous empirical results on the beneficial effect of non-unital noise on the functioning of a gate-based echo state network for temporal processing [31, 46].
Damping prevents the decoherence of reservoir state
Now we proceed by providing such theoretical explanation of the need of damping for the functioning of a gate-based echo state network. A fundamental prerequisite for the effectiveness of a reservoir computer is its capacity to separate distinct inputs reliably. Thus, the reservoir dynamics must be sufficiently expressive to produce distinguishable states $\rho_{x}=\mathcal{U}_{x}\rho$ and $\rho_{y}=\mathcal{U}_{y}\rho$ after input injection. Here, $\mathcal{U}_{j}\,,\,j=x,y$ denotes the unitary encoding of the input value $j$, described in detail in the Methods. The separability ensures that the internal states of the reservoir encode meaningful distinctions in the input, which is essential for downstream tasks such as classification or prediction. As stated by Holevo–Helstrom theorem [47, 48], the distinguishability of two different quantum states may be quantified by their trace distance, namely $\mathcal{T}(\rho,\sigma)=\frac{1}{2}\norm{\rho-\sigma}_{1}$ with $\norm{A}_{1}=\text{Tr}\left(\sqrt{A\,A^{\dagger}}\right)$. By continuity of the unitary evolution, the difference in trace norm of two internal states of the reservoir, corresponding to distinct values of the input, is related to the distance between the initial state $\rho$ and the maximally mixed state ^$\hat{\mathcal{I}}/d$, namely
$\mathcal{T}\left(\rho_{x}-\rho_{y}\right)\leq C(U)\cdot C(|x-y|)\cdot\mathcal{T}\left(\rho-\hat{\mathcal{I}}/d\right)$ (4)
Here, we are denoting with $C(U)$ and $C(|x-y|)$ two constants that depend on the unitary and the distance input values, respectively, and with $d=2^{N}$ the dimension of the Hilbert space. We quantify the expression in Eq. (4) in Fig. 2, showing that a larger deviation from $\hat{\mathcal{I}}$ leads to enhanced separability between distinct inputs. Nonetheless, the repeated mid-circuit measurement leads to quantum decoherence, and the internal state $\rho_{t}$ is driven toward the maximally mixed state, $\rho_{t}\rightarrow\hat{\mathcal{I}}/d$. Indeed, any unitary rotation does not increase the distance from the identity, since
$\norm{\mathcal{U}_{x}\rho-\hat{\mathcal{I}}/d}_{1}=\norm{\mathcal{U}_{x}(\rho-\hat{\mathcal{I}}/d)}_{1}=\norm{\rho-\hat{\mathcal{I}}/d}_{1}$ (5)
while projective measurements decrease it, namely
$\norm{\Pi\rho\Pi^{\dagger}-\hat{\mathcal{I}}/d}_{1}=\norm{\Pi(\rho-\hat{\mathcal{I}}/d)\Pi^{\dagger}}_{1}\leq\norm{\rho-\hat{\mathcal{I}}/d}_{1}\,.$ (6)
Here the equality is realized if and only if $\rho=\bigotimes^{N_{a}}|i\rangle\langle i|$, with $i=0,1$, namely is the tensor product on the eigenvalues of the projective measure. Consequently, the dynamics produces progressively less distinguishable states. Eventually, after an effective time window, the state of the reservoir becomes independent from the input, since $\mathcal{U}_{x}\,\hat{\mathcal{I}}=\hat{\mathcal{I}}$ for any input value $x$. In the following, we demonstrate how dissipation prevents the loss of information, enabling stable computation over an extended period. In particular, the concentration of the reservoir in the maximally mixed state $\hat{\mathcal{I}}$ can be prevented by exploiting the action of a non-unital channel. We recall that a quantum channel $\mathcal{E}$ is non-unital when $\mathcal{E}(\hat{\mathcal{I}})\!=\hat{\mathcal{I}}$. As an example, the amplitude damping channel can be adopted for this purpose, as it tends to relax the qubit state toward |$\ket{0}$, with a decay rate denoted with $\gamma$ throughout the work. Moreover, if each qubit is affected by damping, the effect of dissipation is not hindered by the dimensionality of the circuit, as shown in Fig. 2.
Universal approximation property ensured by non-unital and contractive channel
As a further theoretical support explaining the crucial role of damping, or other similar non-unital channels, we demonstrate that its effect guarantees the universal approximation property of the gate-based echo state network considered in this work and Ref. [31]. A family of computational architectures is universal for a certain set of functionals if, for any of these mappings, there exists an instance within the family that approximates it with arbitrary precision [43]. In particular, the universality of a family of reservoir computers for fading memory mappings relies on some general sufficient conditions [10, 42]. Precisely, a reservoir computer is universal for temporal tasks if the internal dynamics of the reservoir has fading memory and can discriminate between distinct inputs. The formal theorem is detailed in the Supplementary Material attached to this work. Then, the polynomial subalgebra of functionals is constructed by spatial multiplexing, as shown in Ref. [29]. In particular, we recall that a noise modelled by a non-unital and strictly contractive channel ensures the fading memory of the reservoir computer, as already proved in Ref. [31]. Moreover, we prove in the Appendix that even separability is ensured when the quantum channel has a unique fixed point away from the identity. A similar result was stated and proved in Ref. [49, Theorem 2] . This is the case, for example, for the amplitude damping channel. Thus, we prove that a gate-based quantum echo state network is universal under the effect of a class of noise that includes the amplitude damping, thus providing further support to the empirical findings. We refer to the Supplementary Materials for the detailed mathematical proof.
@@
where the vector $\overline{z}^{i}_{t}=\left\{(z_{t})_{s}\right\}_{s=1,\dots,S}$ contains the outcomes of measurements for each repetition of the experiment. Then, we compute the mean correlation over the whole execution by averaging over $t$. Fig. 3 shows the mean correlation over time for increasing time windows $k$. As expected, higher noise intensity boosts the progressive loss of correlation. The loss of correlation appears comparable for different noise models. Thus, fading memory alone does not a priori justify the beneficial effect of amplitude damping on the short-term memory capacity of the network and its accuracy in learning tasks. From a theoretical point of view, fading memory is related to the contractivity of the evolution of the echo state network [44]. In our model, the contractivity in trace norm is ensured by the action of noise, as discussed in the Supplementary Material.
@@
Figure 3: Effect of noises on learning. a) Decay over time in the mean correlation in subsequent measurements is more pronounced when amplitude damping intensity increases. Yellow lines indicate a higher intensity of the noise. The thickness of lines is proportional to the accuracy in NARMA5 prediction. b) From left to right, the blue area is the image of the quantum channel on the Bloch sphere. Black and blue dots are the projections of the sphere’s center, respectively, before and after the application of the channel. For clarity, simulations are for $\gamma=0.6$ and tenfold intensity of noise in the IBM_OSAKA backend. c) Accuracy in NARMA-8 task and short-term memory capacity for various numbers of measured qubits in the 7 qubits register, under the effect of increasing amplitude damping. $\gamma=0$ amounts for noiseless execution of the circuit. Yellow lines correspond to complete measurements. Running average on a window of ten values of $\gamma$ is performed for smoother visualization. d) Accuracy in NARMA-n tasks under the effect of increasing amplitude damping. Here, the reservoir is made of 7 qubits, and all the qubits are measured in the experiment. Results for a swarm of 20 executions are plotted in transparency, mean values are in bold in the picture. Running average on a window of ten values of $\gamma$ is performed for smoother visualization. The optimal noise regime corresponds to $\gamma\sim 0.03$. e) Accuracy in reproducing NARMA tasks and memory capacity for increasing intensity of amplitude damping (in blue), phase damping (in grey), and depolarizing channel (in green). The performance of the noiseless execution corresponds to $\gamma=0$. Only amplitude damping ensures the learning of the network. Again, the complete measurement of a reservoir with 7 qubits is employed
@@
Figure 4: Reconstruction of NARMA tasks on simulators and quantum hardware. a) The reconstruction of NARMAn tasks for, respectively, the noiseless execution of the circuit, the execution affected by optimal amplitude damping, and the realistic noise emulating the one in IBM_OSAKA quantum backend. Optimal values of noise are $\gamma=0.024,0.03,0.014$ for NARMA2, NARMA5 and NARMA8, respectively. The grey vertical line separates the training and test intervals in the output sequence. The amplitude damping ensures the best learning of the system, which is slightly deteriorated under the composition of several noisy quantum channels representing the realistic noise. We highlight that, in the noiseless execution, the target is met for around 50 time-steps. After that, the internal state of the reservoir is driven to the maximally mixed state, becoming unresponsive as discussed in Fig. 2. b) The experimental results of reconstruction of NARMA2, NARMA5 and NARMA8 (from left to right) tasks employing IBM_BRISBANE quantum backend. The experiment is repeated 5 times. The mean results are reported in the green line, with error bars indicating a $1\,\mathrm{s.d.}$ interval. The input length is reduced, in order to reduce the depth of the circuit to fit the technical limitations of the quantum hardware. Precisely, we sample 80 values for $u_{t}$. Reconstruction deteriorates for increasing the depth in memory required, indicating the necessity of the implementation of error mitigation techniques that allow for reducing the unital noise in the evolution.
Effect on learning of simulated quantum noise models
@@
In this Section, we study the effect of different noise models on the memory capacity and the learning capability of a gate-based echo state network [31]. Concerning the simulation of diverse models of quantum noise, their effect on the learning capabilities of the network is investigated by systematically tuning its intensity in numerical experiments. Throughout the whole Results section, the parameter that tunes noise intensity is referred to as $\gamma$. Our results confirm previous findings [31, 39], showing that amplitude damping is drastically improving the performance of the reservoir computer. After confirming that any contractive noisy channel ensures the fading memory of the network, we investigate their positive on the predictive capabilities of the architecture using two standard benchmarks for recurrent networks, namely the measure of short-term memory capacity (MC) and the non-linear autoregressive moving average (NARMA) task, presented in detail in Subsection ”Benchmark for performance analysis”. The short-term memory capacity measures the ability of the network to reproduce online a delayed window of the time-dependent input, while the NARMA task requires approximating a non-linear functional that depends on a fixed amount of the recent information in the input time series. Three noise models are applied separately, namely amplitude damping, phase damping, and depolarizing. Among them, only the amplitude damping (AD) is non-unital. Remarkably, we observe that the action of the only amplitude damping drastically enhances the memory capacity of the network. Consistently, the higher memory capacity is reflected in better accuracy in reproducing the time-dependent task. On the other hand, phase damping (PD) and depolarizing (DE) noise have detrimental effects on both memory capacity and accuracy. These results are summarized in Fig. 3, showing the normalized mean square error (NMSE) and MC for increasing the intensity of the three noises considered. The beneficial effect of amplitude damping is justified by its non-unital nature as a quantum channel. Indeed, the action of a non-unital channel enriches the dynamics of the reservoir allowing for a wider exploration of the phase space and, consequently, greater storage of information. We discuss more details about non-unital channels and theoretical insight into the role of unitality in quantum reservoir computing in the Supplementary Material. To anticipate the behavior on NISQ hardware, we repeated the experiment by applying a channel that replicates the noise model of the backends provided by IBM. A realistic noise model for current superconducting hardware consists of the application of the composition of quantum channels, including, but not restricted to, the three noise models considered in this work. Thus, recalling that quantum channels are non-expanding by definition [49], we may assume that the dynamics is strictly contractive since the composition of non-expanding channels and strictly contractive channels is in turn a contractive channel. Moreover, we confirmed that the realistic noise is non-unital by analyzing its effect on the Bloch sphere, as shown in Fig. 3. As expected, the network still shows a remarkable improvement in learning capabilities when compared with noiseless executions, even if the accuracy in reproducing tasks is slightly deteriorated compared to optimal pure amplitude damping. The results in the reconstruction of the NARMA task are reported in Fig. 4.